OpenStudy (anonymous):

sqrt(-16)= ? simplify and express in terms of i

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

in terms of i........would that be 4i ?

OpenStudy (anonymous):

Yep, = 4i

OpenStudy (anonymous):

ok, so 4i and not -4i?

OpenStudy (anonymous):

√(-16) = √(-1) * √16 = i * 4 = 4i

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

nope no more confusion :)

OpenStudy (anonymous):

:)